Solve the system by addition. 5x – 3y = 13
4x – 3y = 11

Answer by jai_kos(139)

Given 5x – 3y = 13
4x – 3y = 11
Subtract both equations, we get
5x - 4x = 13 -11
x = 2
Substitute the value of x = 2 in the equation (1), we get
4(2)–3y = 11
8–3y =11
Get all the numbers to one side, we get
8 - 11 = 3 y

- 3 = 3y
Divide by 3 , we get
-1 = y
Hence the value of “x” =2 and y = -1